OpenStudy (anonymous):

what is 1 1/3 x 3 3/4 ? (how to solve?)

OpenStudy (anonymous):

1 1/3 = 4/3 3 3/4 = 15/4 4/3 * 15/4 = 5

OpenStudy (anonymous):

remember 1 1/3 is the same thing as 1+1/3 and 3 3/4 is the same thing as 3+3/4.